# Large-time Behavior for a Cutoff Level-Set Mean Curvature G-equation with Source term

### v1

@@ 1. Introduction and Outline of Paper
where $g,f\in C(\mathbb{R}^{n})$ are a given initial datum and non-negative source, respectively. The cutoff function, which is the source of full nonlinearity, is defined by $(\cdot)_{+}:=\max\{0,\cdot\}$. When describing front propagation, the cutoff is a mathematical tool that prohibits any retreating movement of the evolving front. We also introduce a wind-type vector field $\vec{W}$, which can be an incompressible mixing flow. This term appears in the so-called G-equation, which is used in premixed turbulent combustion applications. (C) not only has motivations arising from turbulent combustion theory, but also modeling of crystal growth. See below in the Motivations subsection.
The well-posedness theory of (C) is already established, see [2, 3, 4, 6]. We will consider finer properties, such as large-time behavior and regularity, under the viscosity solution framework. We emphasize that the structure of the nonlinearity is non-coercive with respect to the gradient term, which creates some challenges in establishing a priori gradient estimates. Moreover, there is no convexity with respect to the gradient term, which does not allow for the use of standard large-time behavior results in the Hamilton-Jacobi equation theory. Despite these challenges, we have monotonicity of the nonlinear term, namely its non-negative definiteness. Using this structure, we are able to study these deeper properties of viscosity solutions. In this work, we focus on studying (C) in two settings: periodic and radially symmetric.
@@ 1.1. Motivations
From the combustion theory/literature an interesting model arises known as a cutoff G-equation with a shear-flow structure. This model is used to describe the motion of a flame front under some wind with the physical restriction that once a region has been burned, it cannot be burned again. This behavior is apparent in wild forest fires, for example. The flame front propagation can be described by the (n1)(n-1)-dimensional minimal surface-type equation
@@
see [15]. We can think of (C) as the $n$-dimensional non-graph mean curvature flow of the flame propagation with the cutoff function prohibiting non-physical burning of already burned regions. Studying the periodic case for the stationary problem corresponds to the ergodic cell problem in homogenization and large-time dynamics theory. This model has been studied by the authors of [15]. In particular, they have proved the existence of an effective burning velocity for a range of flow intensities. This is also known as the large-time average. In fact, they showed that if the intensity of the shear-flow is greater than a critical value, there is no asymptotic speed for the evolving flame front, and therefore no hope for convergence to an asymptotic profile. See Section $1$ of [15] for the main results regarding this. The question whether there is convergence in the suitable range of flow intensities is still open. However, this paper can provide a positive result if a source term appears in the equation. For more works regarding turbulent combustion as well as a two-player game approach, we refer the reader to [5, 10, 20].
On the other hand, (C) has motivations arising from crystal growth applications. For example, the following birth and spread model was introduced in [8]:
@@ 2. Periodic setting
with $a^{ij}(Du):=\left(\delta_{ij}-\frac{u_{x_{i}}u_{x_{j}}}{|Du|^{2}}\right)$ and cutoff $(\cdot)_{+}:=\max\{0,\cdot\}$. Here, $\delta_{ij}$ is the Kronecker delta. Moreover, $g\in C(\mathbb{T}^{n})$ is a given initial datum and $f\in C(\mathbb{T}^{n})$ is a given non-negative source. The so-called wind flow $\vec{W}$ is a Lipschitz vector field. Lastly, we define $\mathbb{T}^{n}:=\mathbb{R}^{n}/\mathbb{Z}^{n}$ as the usual $n$-dimensional flat torus. We will also consider the corresponding ergodic cell problem
@@ 4.1. Local Lipschitz Regularity in the Periodic Setting
Therefore, $u$ is Lipschitz in time at $t=0$. Now fix some $s>0$, and let $v(x,t):=u(x,t+s)$. Since the PDE is invariant in time, $v$ is another viscosity solution to (N) with different initial datum $v_{0}(x):=v(x,0)=u(x,s)$. Denote $u_{0}(x):=u(x,0)$. Observe that

### v2

@@ 1. Introduction and Outline of Paper
where $g,f\in C(\mathbb{R}^{n})$ are a given initial data and non-negative source, respectively. The cutoff function, which is the source of full nonlinearity, is defined by $(\cdot)_{+}:=\max\{0,\cdot\}$. When describing front propagation, the cutoff operator is a mathematical tool that prohibits any retreating movement of the evolving front. We also introduce a wind-type vector field $\vec{W}$, which can be an incompressible mixing flow. This term appears in the so-called G-equation, which is used in premixed turbulent combustion applications [19, 22]. Equation (C) not only has motivations arising from turbulent combustion theory, but also from modeling of crystal growth. See below in the Motivations subsection. Lastly, when describing front propagation, equation (C) can be considered as a level set equation. That is, the zero-level set of the solution corresponds to the hypersurface of the front. See [17, 18] for more information on the level set method.
The well-posedness theory of (C) is already established, see [2, 3, 4, 6]. We will consider finer properties, such as large-time behavior and regularity, under the viscosity solution framework. Such studies for similar first-order and second-order Hamilton-Jacobi equations with convex and coercive Hamiltonians have been completed in the past three decades. See [16], for example. We emphasize that the structure of the nonlinearity in (C) is non-coercive with respect to the gradient term, which creates some challenges in establishing a priori gradient estimates. Moreover, there is no convexity with respect to the gradient term, which does not allow for the use of standard large-time behavior results in the Hamilton-Jacobi equation theory. Despite these challenges, we have monotonicity of the nonlinear term, namely its non-negative definiteness. Using this structure, we are able to study these deeper properties of viscosity solutions. In this work, we focus on studying (C) in two settings: periodic and radially symmetric.
@@ 1.1. Motivations
From combustion theory, an interesting model arises: the cutoff G-equation with a shear-flow structure, [13, 14, 15]. This model is used to describe the motion of a flame front under some wind with the physical restriction that once a region has been burned, it cannot be burned again. This behavior is apparent in wild forest fires, for example. The flame front propagation can be described by the (n1)(n-1)-dimensional minimal surface-type equation
@@
see [15]. We can think of (C) as the $n$-dimensional non-graph mean curvature flow of the flame propagation with the cutoff function prohibiting non-physical burning of already burned regions. Studying the periodic case for the stationary problem corresponds to the ergodic cell problem in homogenization and large-time dynamics theory. Finer properties of the solutions to this model have been studied by the authors of [15]. In particular, they have proved the existence of an effective burning velocity for a range of flow intensities. This is also known as the large-time average. In fact, they showed that if the intensity of the shear-flow is greater than a critical value, there is no asymptotic speed for the evolving flame front, and therefore no hope for convergence to an asymptotic profile. See Section $1$ of [15] for the main results regarding this. The question whether there is convergence in the suitable range of flow intensities is still open. However, this paper can provide a positive result if a source term appears in the equation. For more works regarding turbulent combustion as well as a two-player game approach, we refer the reader to [5, 11, 20, 23]. For a periodic homogenization result for this model, see [10]. Lastly, see [24] for simulations of front propagations.
On the other hand, (C) has motivations arising from crystal growth applications. For example, the following birth and spread model was introduced in [7] and [8]:
@@ 2. Periodic setting
with $a^{ij}(Du):=\left(\delta_{ij}-\frac{u_{x_{i}}u_{x_{j}}}{|Du|^{2}}\right)$ and cutoff $(\cdot)_{+}:=\max\{0,\cdot\}$. Here, $\delta_{ij}$ is the Kronecker delta. Moreover, $g\in C(\mathbb{T}^{n})$ is a given initial data and $f\in C(\mathbb{T}^{n})$ is a given non-negative source. The so-called wind flow $\vec{W}$ is a Lipschitz vector field. Lastly, we define $\mathbb{T}^{n}:=\mathbb{R}^{n}/\mathbb{Z}^{n}$ as the usual $n$-dimensional flat torus. We will also consider the corresponding ergodic cell problem
@@ 4.1. Local Lipschitz Regularity in the Periodic Setting
Therefore, $u$ is Lipschitz in time at $t=0$. Now fix some $s>0$, and let $v(x,t):=u(x,t+s)$. Since the PDE is invariant in time, $v$ is another viscosity solution to (N) with different initial data $v_{0}(x):=v(x,0)=u(x,s)$. Denote $u_{0}(x):=u(x,0)$. Observe that
@@ 4.2. Lipschitz Regularity in Radially Symmetric Setting
Acknowledgments
The author would like to thank Professors Hung V. Tran and Hiroyoshi Mitake for their invaluable feedback and suggestions. Moreover, the comments provided by Jiwoong Jang are greatly appreciated.